solve:
-3(4.2-3.5x=8x-7.2
HOW DO YOU SOLVE THIS PROBLEM!??!

Answer by Alan3354(69443)

solve:
-3(4.2-3.5x=8x-7.2
----------------
There's a right parenthesis missing.
-3(4.2-3.5x)=8x-7.2 I added one. If it's in the right place, the answer will be correct.
-3(4.2-3.5x)=8x-7.2
expand things
-12.6 + 10.5x = 8x - 7.2
Subtract 8x
-12.6 + 2.5x = -7.2
Add 12.6
2.5x = 5.4
x = 5.4/2.5
x = 2.16